OpenStudy (anonymous):

Help please! If (x + k) is a factor of f(x), which of the following must be true? f(k) = 0 f(–k) = 0 A root of f(x) is x = k. A y intercept of f(x) is x = –k.

OpenStudy (irishboy123):

If (x + k) is a factor of f(x), then f(x) = (x+k).g(x) i think you can work it out from there

OpenStudy (anonymous):

C??? because if f(x)=x+k) then wouldnt a root of f(x) x=k

OpenStudy (dan815):

f(-k)=0

OpenStudy (phi):

almost by definition of factor f(x)= (x+k)* g(x) what do we get when x= -k ? putting in -k for x we get f(-k) = (-k+k) * g(-k) we don't know what g(-k) is, but we do know what (-k+k) is what do we get for f(-k) ?

OpenStudy (anonymous):

the answer is A

OpenStudy (phi):

do you see f(-k) = (-k+k) * g(-k) becomes f(-k) = 0 * g(-k) and 0 * anything is 0